# SecBench: A Comprehensive Multi-Dimensional Benchmarking Dataset for LLMs in Cybersecurity

### v1

@@ 5.1. MCQ Benchmarking
Qwen2-7B(Qwe, 2024) 87.74 88.29 82.29 90.77 75.29 86.94 85.79 85.20 89.38 90.69 83.41 82.49 91.86 83.74
Yi-1.5-9B(Yi_, 2024) 86.44 87.03 80.57 89.04 75.74 85.58 85.61 83.90 87.19 89.74 83.93 80.76 89.63 82.22
Llama-3-70B(lla, 2024) 88.86 89.46 82.97 90.95 80.28 87.27 87.95 85.95 88.54 92.81 87.96 81.95 92.44 87.24
Qwen2-72B(Qwe, 2024) 92.41 92.71 89.50 94.50 83.83 91.90 91.55 91.04 93.01 94.56 90.78 90.05 94.26 89.13
Yi-1.5-34B(Yi_, 2024) 89.59 90.04 85.19 91.48 81.82 89.14 88.71 88.47 90.00 92.20 88.44 87.14 91.38 84.15
@@
Qwen2-7B(Qwe, 2024) 59.99 53.39 63.79 60.14 54.25 63.67 64.71 53.67 62.61 57.75 60.80 55.94 63.78 67.57
Yi-1.5-9B(Yi_, 2024) 65.24 64.98 65.39 65.01 73.88 65.86 67.04 62.61 63.77 64.47 66.57 66.88 66.36 74.23
Llama-3-70B(lla, 2024) 68.12 65.93 69.47 69.48 20.62 71.20 67.35 65.81 70.87 67.64 72.67 71.61 67.31 63.14
Qwen2-72B(Qwe, 2024) 82.13 75.93 85.70 81.96 88.25 84.82 85.45 76.89 86.85 80.52 83.21 80.00 85.14 86.13
Yi-1.5-34B(Yi_, 2024) 75.03 67.82 79.18 74.75 85.38 79.25 78.27 68.05 76.71 73.71 77.37 62.50 79.25 81.80

### v2

@@ 5.1. MCQ Benchmarking
Qwen2-7B-Instruct(Qwe, 2024) 87.74 88.29 82.29 90.77 75.29 86.94 85.79 85.20 89.38 90.69 83.41 82.49 91.86 83.74
Yi-1.5-9B-Chat(Yi_, 2024) 86.44 87.03 80.57 89.04 75.74 85.58 85.61 83.90 87.19 89.74 83.93 80.76 89.63 82.22
Llama-3-70B-Instruct(lla, 2024) 88.86 89.46 82.97 90.95 80.28 87.27 87.95 85.95 88.54 92.81 87.96 81.95 92.44 87.24
Qwen2-72B-Instruct(Qwe, 2024) 92.41 92.71 89.50 94.50 83.83 91.90 91.55 91.04 93.01 94.56 90.78 90.05 94.26 89.13
Yi-1.5-34B-Chat(Yi_, 2024) 89.59 90.04 85.19 91.48 81.82 89.14 88.71 88.47 90.00 92.20 88.44 87.14 91.38 84.15
@@
Qwen2-7B-Instruct(Qwe, 2024) 59.99 53.39 63.79 60.14 54.25 63.67 64.71 53.67 62.61 57.75 60.80 55.94 63.78 67.57
Yi-1.5-9B-Chat(Yi_, 2024) 65.24 64.98 65.39 65.01 73.88 65.86 67.04 62.61 63.77 64.47 66.57 66.88 66.36 74.23
Llama-3-70B-Instruct(lla, 2024) 68.12 65.93 69.47 69.48 20.62 71.20 67.35 65.81 70.87 67.64 72.67 71.61 67.31 63.14
Qwen2-72B-Instruct(Qwe, 2024) 82.13 75.93 85.70 81.96 88.25 84.82 85.45 76.89 86.85 80.52 83.21 80.00 85.14 86.13
Yi-1.5-34B-Chat(Yi_, 2024) 75.03 67.82 79.18 74.75 85.38 79.25 78.27 68.05 76.71 73.71 77.37 62.50 79.25 81.80